OpenStudy (dtan5457):

The heights, h, of the students in the chorus at Central Middle School satisfy the inequality (below), when h is measured in inches. Determine the interval in which these heights lie and express your answer to the nearest tenth of a foot.

OpenStudy (dtan5457):

\[\left| \frac{ h-57.5 }{ 2 } \right|\le 3.25\]

OpenStudy (dtan5457):

@Nnesha @KamiBug @freckles

OpenStudy (dtan5457):

Nvm I just answered my own question..

Nnesha (nnesha):

good :D::D:D: